# From Cash to Cashless: UPI’s Impact on Spending Behavior among Indian Users

### v1

@@ 4.2.1. EXPLORING UPI USAGE FOR TRANSACTIONS: INSIGHTS AND PERSPECTIVES
Respondents shared valuable insights regarding their experiences with UPI usage in our research findings, revealing common themes, patterns, and occasional divergent viewpoints. Many respondents lauded UPI for its convenience as a payment method, emphasizing its ability to eliminate the need to carry money physically. This convenience streamlined small payments, rendering them more efficient. Moreover, the consistent praise for UPI transactions centered on their remarkable speed and efficiency. Respondents frequently described these transactions as exceedingly fast and convenient.
@@
According to a State Bank of India report, for every Rs 1 (0.012 USD) increase in UPI, there is an 18 paisa (0.0022 USD) decline in debit card transactions, indicating that the number of ATM visits has decreased from 16 to 8 times a year (dho, 2023). Also, the fact the transaction creates a digital record helps them keep a record of transactions and financial documentation. These factors collectively contribute to the high satisfaction levels among UPI users
@@
“I love UPI! It’s a fantastic way to spend money without dealing with change or carrying cash. You can make payments from just Rs 1 (0.012 USD) to a whopping 1 lakh rupees (12,028.79 USD) in one go. It’s one of the best things in this progressive era.”-[P1]
The cashless nature of UPI was another aspect that resonated with several participants, simplifying the payment process by obviating the requirement to carry tangible currency. Furthermore, certain respondents, particularly a lawyer and a clinic owner, emphasized how UPI had markedly enhanced their business operations by expediting payments. This improvement in efficiency contributed to a smoother and more streamlined financial process for these businesses.
“As a lawyer, frequent photocopies and official expenses often require change money. However, I typically only carry a little cash with me. This led to a time-consuming process, motivating me to embrace UPI for seamless small payments.”-[P2]
One salient feature of UPI is that it has no processing/transaction fees (for P2P transactions), unlike credit and debit cards. Also, processing fees may or may not be there, depending upon the transaction amount and other factors. This benefits not only the merchants as well as the customer.
@@
”I encourage my customers to use UPI or cash for payments instead of credit or debit cards to avoid the 2-3% transaction fees, which either come from my earnings or are passed on to the customer, potentially harming my reputation. UPI offers a real-time, no transaction fees and transparent payment solution.”-[P3]
Many respondents believe that possessing financial proficiency is essential when utilizing UPI. Individuals must delve into personal finance education and exercise restraint in expenditures to mitigate these issues. Improving money management habits is imperative to maintain financial stability while utilizing UPI.
”Personally, there are occasions when I exceed my initial spending plans, but I rely on finance tracking applications to ensure I stay within my budget.”-[P4]
The adoption of UPI has not been without its share of risks and challenges.A recurring theme was the awareness of potential for scams and fraudulent activities and its rising numbers. Several respondents recounted incidents involving phishing attempts, where malicious actors impersonated their known ones or UPI platforms to deceive users into sharing sensitive information. Additionally, users recounted instances of phishing messages disguised as official UPI notifications, urging them to click on malicious links and disclose sensitive information such as OTPs (One-Time Passwords) and bank details. These fraudulent activities have raised questions about the vulnerability of users to cyber threats. To address these challenges, users can take proactive measures to protect themselves. Recommendations based on user experiences include avoiding sharing sensitive information, such as OTPs, with anyone, regularly updating UPI apps, enabling two-factor authentication, and being cautious when receiving unsolicited payment requests. According to the Ministry of Finance, more than 95,000 fraud cases linked to UPI transactions were recorded in 2022-23, an increase from 84,000 incidents in 2021-22 (sin, 2023).
@@
“I once had a bad experience with UPI. I got a message asking me to send money, and they wanted my bank OTP. Sadly, I gave it to them, and things got messed up. They took a big chunk of cash, around 10,000 rupees, from my bank account. Fortunately, my father worked at a bank and helped fix it. I learned a harsh lesson about being careful with UPI and not sharing important info.”-[P6]
”Whether UPI is good depends on how careful you are with it and how well you handle your money. If you’re smart about it, it can be useful. For example, my friend helped me get my money back from a scammer using UPI, which was great. But you’ve got to be cautious and use it wisely, or it might not turn out well.”-[P5]
@@ 4.2.2. FACTORS INFLUENCING SPENDING BEHAVIOR
A few respondents strongly preferred using UPI as a payment option. It simplifies transactions, reduces dependency on cash, and encourages frequent expenditure. This shift contributes to a more efficient, cashless economy, revolutionizing how people handle their finances and everyday purchases.
”During and after COVID, more transactions moved online, making things even more convenient. This helped the economy grow. When I have less cash, I tend to spend less. But I spend more often with UPI because there’s no need to hold back. You don’t have to keep track of physical cash; you can check your bank balance and pay directly from your bank account.”-[P7]
@@
“UPI has made transactions easier. Combined with the rise of Swiggy/Zomato/Zepto/Blinkit, etc., ordering online has become much more appealing and easy. Without UPI, I might not have been spending much on food online. It’s so easy; no need to think twice. Sometimes, I regret purchases because you only need a phone with wifi/mobile data”-[P8]
This digital payment method has reshaped how individuals perceive and manage their finances, sometimes leading to more frequent and impulsive spending patterns.
@@
”Cash feels like something tangible, like giving away something valuable when we spend it. But with UPI, it’s just a number on the screen, and it doesn’t feel the same way.”-[P9]
“Running out of cash is never a problem now. Spending does increase as you do not feel the pinch of physical cash being spent. As you always carry your whole bank balance, spending increases.”-[P10]
“Digital money does not give you guilt when spending. When you spend paper money, you feel guilty.”-[P11]
Many people often struggle with regret over their financial choices, and the reasons behind this regret are complex. This regret often comes from spending too much, needing to be mindful of their expenses, and making less-than-ideal financial decisions. The issue’s core lies in the sharp difference between physical cash and the abstract nature of digital transactions (UPI payments). While cash represents a physical object that one can touch and watch disappear from their wallet, digital transactions involve only numbers on a screen, disconnected from the physical world. This lack of a tangible connection and the absence of a solid feeling about the money being spent lead to spending without much thought, creating a sense of emptiness later filled with regret and guilt.
@@
When we asked users about their spending habits and where they generally use UPI, a diverse range of insights emerged. It was evident that UPI has found its way into various aspects of daily life, transforming how people manage their expenses. Users reported using UPI for essentials like rent, groceries, and bills, as well as for work-related expenses, leisure activities such as dining out and entertainment, education, online shopping, and even small everyday indulgences like street food and auto-rickshaw fares.
I have increased expenditure on minor items such as tea or lime soda! However, since most establishments in my area now support UPI payments, it’s convenient when I forget to carry cash. I seldom have any physical currency with me these days. -[P12]
Additionally, some users rely on UPI for critical business transactions, emphasizing its role in personal and professional financial activities. With the integration of UPI as a mode of payment in almost every app, UPI is ubiquitous and a definite game changer in the Indian economic landscape.
Most respondents agreed with the statement that using UPI affects their ability to stick to their budget. They either agreed or strongly agreed that UPI impacts their budgeting habits. This suggests that for many users, UPI does influence their spending behavior and financial discipline.
“UPI brought the revolutionary change in the payment system but also introduced the bad habits of overspending money. But, be careful when it starts allowing credit payments, as that might tempt people to spend more.”-[P13]
Our study uncovers clear trends in UPI adoption across user groups. Working professionals and business owners have seamlessly integrated UPI into their financial routines, using it for bill payments, invoicing clients, and managing business expenses. On the other hand, students from various age groups are adopting UPI for academic needs, such as paying tuition fees, buying books, and subscribing to online courses. Common spending trends, such as grocery shopping, cab bookings, and food delivery, were observed consistently across age groups and professions, demonstrating the broad utility of UPI for routine transactions and its transformative influence on spending behavior.
@@ 4.2.3. ATTITUDE TOWARDS FINANCIAL MANAGEMENT
Respondents held diverse perspectives on their financial discipline, with some expressing confidence in their ability to manage finances (rating 4 or 5) while others appeared less assured (giving ratings of 2 or 3). This shows that people have different levels of comfort with money matters.
Opinions about how UPI affects financial management were mixed. Some found it helpful, while others felt it complicated things. This suggests that the impact of UPI on personal finances varies depending on individual preferences and habits.
“Since I can refer to my balance and track my spending, controlling my expenses and budget has become easier. The same may not be the case for many of the population since the accessibility feature of UPI makes it harder for people to stick to a budget, which leads to overspending.”-[P14]
@@ 4.2.4. RECOMMENDATIONS FOR FINANCE TRACKING APPLICATIONS
“I’d like my UPI app to have some specific features. It should allow me to restrict spending in certain categories, such as food, by using QR codes to track locations and deny transactions in those places. I want the app to generate a pie chart showing the average payments I make in different categories over a specific period, making tracking my expenses easier.”-[P15]
Responses from various professional backgrounds revealed that the specific needs of their professions often shape users’ expectations of financial management apps. Business professionals, including entrepreneurs, sought tools to manage business-related finances efficiently, such as tracking client payments and categorizing expenses. Lawyers expressed the need for streamlined record-keeping for legal practices, while doctors aimed to collect patient payments and related transactions effectively. Working-class professionals valued simplicity and ease of use in financial apps, emphasizing graphical representations and clear spending patterns. These insights underscore the importance of tailoring financial apps to meet the unique requirements of diverse professions, ensuring effective financial management across various occupational backgrounds.
“As a doctor, it would be beneficial if I could easily keep track of all the patients who have paid their fees. Checking transaction histories repeatedly can be quite a hassle. So, having an automatically generated list of patients who have made payments, with a clear categorization of whether the payment has been completed, would be incredibly convenient.”-[P16]
Some respondents went beyond traditional budgeting and desired financial education and guidance within the app. They emphasized the importance of learning to achieve financial stability and discipline through the application’s features.
“Besides handling expenses and transactions, it would be great if the app also offers guidance on achieving financial stability and discipline, perhaps by providing tips and educational content to help users make better financial decisions.”-[P17]
Respondents highlighted the importance of receiving notifications and alerts when approaching spending limits or exceeding budgeted amounts. This feature is a valuable tool to maintain financial discipline.
“An app where we can set limits for miscellaneous grocery shopping expenses and alert users about their financial condition. I’d appreciate having a daily spending limit feature that sends me reminders when I approach or reach my daily spending goals, helping me manage my finances more effectively.”-[P18]
@@ 5. Discussion
FEATURES
YEARLY SUBSCRIPTION COST
PROS
CONS
SUPPORTING PLATFORMS
FinArt (fin, 2016)
Automatically organizes transactional SMS into expenses, budget, and bill reminders, even when the app is closed.
Rs 799 (9.61 USD)
Data stored on Drive, Auto Tracking, spend analysis dashboards, and budget setting. Separates personal and business accounts for comprehensive financial management.
Setup complexity, ads persist after subscription, and occasional difficulty in SMS tracking.
ANDROID, PC, IOS
Money Manager (mm, 2020)
The app efficiently handles dual accounts, ensuring separate tracking of income and expenses. It offers offline/cloud backup and Data Export functionalities.
Free
Diverse payment method profiles, encompassing cash and cards, support competitive analysis. Easily create new profiles based on specific needs.
The application takes a long time if currencies are changed to optimize expenses and income. Most fees are non-editable.
ANDROID, IOS
Monefy (my, 2016)
The app is straightforward; the infographics make it vibrant and easy to use.
Rs 180 (2.17 USD)
Set bill reminders, manage multiple accounts and easily import data through Excel.
Frequent prompts for premium upgrades
ANDROID, PC ,IOS
Wallet (wal, 2014)
Users can use it for different currency rates worldwide. The Group Sharing option is also available in the application
Rs 700 (8.42 USD)
Monitor group expenses in the totals tab, track recent costs in the activity tab, and official API support available.
Accounts auto-disconnect; unable to modify the initial login currency.
ANDROID, IOS
Splitwise (spl, 2011)
App simplifies bill splitting , not for personal expense tracking. Form groups, manage shared bills, and track dues effortlessly.
Rs 499 (6 USD)
Clean UI. Manages multiple wallets, links bank accounts or Crypto wallets directly, and tracks transactions seamlessly (premium feature).
Tracking and analyzing personal finances can be challenging.
ANDROID, PC, IOS
Spendee (spe, 2013)
The app provides user-friendly expense and income tracking with easy-to-use tags for sorting costs efficiently.
Rs 900 (10.43 USD)
Effortlessly track expenses, share budgets, view spending analytics, and easily monitor debt progress.
Certain features are restricted behind a paywall.
ANDROID, PC, IOS
GoodBudget (gb, 2009)
Budget management application. The app lets you create and refill multiple budgets (envelopes) at intervals like monthly or weekly.
Rs 5817 (70 USD)
Create multiple wallets, set up recurring payments, track credit card and debt. Link bank accounts and enable group sharing (premium only).
Simple UI. Adding transactions is tedious, and there’s no API support.
ANDROID, PC, IOS
Table 1. Comparison of Personal Finance Tracking Apps
@@
Answering RQ1 as mentioned in Section 1, the study results reveal a significant impact of UPI on Indian users’ spending habits. Most respondents (74.2%) accepted that UPI adoption led to increased spending, highlighting the UPI influence on user behavior. The data also indicates that 12% of respondents started using UPI after the demonetization in 2016 and , a 16% and 17.6% in 2020 and 2021 respectively, a period marked by significant economic and social changes. This suggests that external events, such as demonetization and pandemic-related lockdowns, may have been one of the causes of adoption of UPI, influencing spending patterns. Furthermore, 59.8% of respondents acknowledged exceeding their budget or overspending due to UPI usage. The qualitative analysis supports these findings. Overall, the evidence suggests that UPI has contributed to increased spending for most users. Users have increasingly turned to UPI for a wide range of transactions, including P2P (peer-to-peer) and P2M (peer-to-merchant) transactions, with P2P transactions taking the lead (met, 2023). With the integration of UPI as a payment method in multiple apps and places like street vendors, users are more likely to engage in impulse purchases, smaller transactions, and online shopping. Additionally, the qualitative analysis highlighted the psychological shift users experience, with the abstract nature of digital transactions (”just a number on the screen”) leading to impulsive spending. In some cases, regret and guilt hit people after they realize their spending by checking their bank balances or statements, compared to spending physical cash, highlighting the lack of tangibility associated with UPI payments. Answering RQ2 as mentioned in Section 1, user satisfaction levels with 91.5% of the survey respondents reporting satisfaction with the usage of UPI and 95.2% of the survey respondents found making payments via UPI convenient. The qualitative analysis underscores the convenience of UPI transactions, with users praising its speed, efficiency, and the elimination of the need for physical cash. The absence of processing fees, as opposed to credit and debit cards, is also highlighted as a positive factor contributing to user satisfaction. The research findings indicate that users perceive UPI as a secure and efficient payment method, eliminating concerns about carrying large sums of money. This positive perception of convenience and efficiency likely contributes to the widespread adoption and satisfaction with UPI, ultimately shaping users’ preferences for cashless transactions and influencing spending behavior. The research also notes that the cashless nature of UPI enhances business operations for professionals like lawyers, emphasizing the efficiency gained through expediting payments.
@@
Answering RQ3 as mentioned in Section 1, the study identifies variations in UPI adoption across demographic groups, particularly in professions. Working professionals and business owners seamlessly integrate UPI into their financial routines, using it for various purposes such as bill payments and managing business expenses. On the other hand, students adopt UPI for academic needs like paying tuition fees. These differences highlight the importance of tailoring financial solutions to meet the specific needs of diverse user segments. Despite variations in adoption, common spending trends are observed across age groups and professions, demonstrating the broad utility of UPI. The findings suggest that the adoption of UPI is influenced by users’ professional backgrounds and financial needs, showcasing the platform’s adaptability to various user requirements.
5.1. RECOMMENDATION AND DESIRED FEATURE IN FINANCE TRACKING APPLICATION AND UPI
@@ 5.2. FINANCIAL APP ANALYSIS IN INDIA
Stakeholders, including app developers, government agencies, and users, stand to benefit from incorporating our suggestions into future financial tracking app developments. This comprehensive analysis addresses user needs and provides valuable insights for enhancing the overall financial management ecosystem, ultimately advancing effective and user-centric solutions.

### v2

@@ 4.2.1. EXPLORING UPI USAGE FOR TRANSACTIONS: INSIGHTS AND PERSPECTIVES
Respondents appreciated UPI’s convenience and cashless transactions, obviating the requirement to carry tangible currency. They consistently praised the system’s remarkable speed and efficiency in handling small transactions.
@@
According to a State Bank of India (SBI) report reveals that with every Rs 1 (0.012 USD) increase in UPI transactions, there is an 18 paisa (0.0022 USD) decline in debit card transactions, leading to a decrease in ATM visits from 16 to 8 times a year [44]. Additionally, UPI transactions create a digital record that helps users maintain a record of transactions and financial documentation, collectively contributing to high satisfaction levels among UPI users.
@@
“I love UPI! It’s a fantastic way to spend money without dealing with change or carrying cash. You can pay just Rs 1 (0.012 USD) to a whopping 1 lakh rupees (12,028.79 USD) in one go. It’s one of the best things in this progressive era.”-[P1]
Certain respondents, like a lawyer and clinic owner, emphasized UPI’s role in expediting payments and enhancing business efficiency and financial processes. Notably, UPI’s absence of processing fees for peer-to-peer transactions stands out, benefiting merchants and customers and distinguishing it from credit and debit cards.
@@
”I encourage my customers to use UPI or cash for payments instead of credit or debit cards to avoid the 2-3% transaction fees, which either come from my earnings or are passed on to the customer, potentially harming my reputation. UPI offers a real-time, no transaction fees and transparent payment solution.”-[P2]
The rise of UPI adoption brings risks, with a surge in scams and fraud. Respondents shared phishing incidents, where malicious actors posed as known contacts or UPI platforms to trick users into sharing sensitive details raising concerns about user vulnerability to cyber threats. Notably, in 2022-23, over 95,000 fraud cases related to UPI transactions were recorded, a notable increase from 84,000 incidents in 2021-22 (sin, 2023).To address these issues, users can take protective steps like avoiding sharing sensitive information like OTPs and bank details, updating UPI apps, enabling two-factor authentication, and being cautious with unsolicited payment requests.
@@
“I once got a message asking me to send money, and they wanted my bank OTP. Sadly, I gave it to them, and things got messed up. They took a big chunk of cash, around Rs 10,000, from my bank account. Fortunately, my father worked at a bank and helped fix it. I learned a harsh lesson about being careful with UPI and not sharing important info.”-[P3]
@@ 4.2.2. FACTORS INFLUENCING SPENDING BEHAVIOR
A few respondents strongly preferred using UPI as a payment option as it simplifies transactions, reduces dependency on cash, and encourages frequent expenditure. This shift contributes to a more efficient, cashless economy, revolutionizing how people handle their finances and everyday purchases.
@@
”I spend more often with UPI because you pay directly from your bank account. UPI combined with the rise of Swiggy or Zomato or Zepto or Blinkit, etc., ordering online has become much more appealing and easy. Without UPI, I might not have been spending much on food online. Sometimes, I regret purchases because you only needs a phone with Wifi/mobile data”-[P4 & P5]
This digital payment method has reshaped how individuals perceive and manage their finances, sometimes leading to more frequent and impulsive spending patterns. Many individuals struggle with financial regret from overspending, the need for expense mindfulness, and less-than-ideal financial choices. The root cause lies in the stark contrast between physical cash and the abstract nature of digital transactions like UPI payments. Physical cash provides a tangible connection, whereas digital transactions involve mere numbers on a screen, leading to spending without much thought and subsequent feelings of regret and guilt.
@@
“Digital money does not give you guilt when spending. When you spend paper money, you feel guilty. Cash feels tangible, like giving away something valuable when we spend it. But with UPI, it’s just a number on the screen, and it doesn’t feel the same way.”-[P6 & P7]
@@
A diverse range of insights emerged when we asked users about their spending habits and where they generally used UPI. It was evident that UPI has found its way into various aspects of daily life, transforming how people manage their expenses. Users reported using UPI for essentials like rent, groceries, and bills, as well as for work-related expenses, leisure activities such as dining out and entertainment, education, online shopping, and even small everyday indulgences like street food and auto-rickshaw fares.
UPI’s integration into numerous applications has positioned it as a key player in personal and professional finances, particularly in critical business transactions. Its omnipresence showcases its transformative impact on the Indian economic landscape. Respondents acknowledge UPI’s significant influence on spending behavior, budgeting habits, and financial discipline. The study identifies distinct and common patterns in UPI adoption among various user groups, emphasizing its broad utility in routine transactions and reshaping spending habits.
@@ 4.2.3. ATTITUDE TOWARDS FINANCIAL MANAGEMENT
Respondents highlight the significance of financial proficiency in UPI usage, stressing the need for personal finance education and expenditure restraint. Attitudes toward financial discipline varied, reflecting different comfort levels in managing finances using UPI, indicating individual variability in its impact on personal finances.
”Personally, there are occasions when I exceed my initial spending plans, but I rely on finance tracking applications to ensure I stay within my budget.”-[P8]
@@ 4.2.4. RECOMMENDATIONS FOR FINANCE TRACKING APPLICATIONS
“I’d like my UPI app to allow me to restrict spending in certain categories, such as food, by using QR codes to track locations and deny transactions in those places. I want the app to generate a pie chart showing the average payments I make in different categories over a specific period.”-[P9]
Responses from various professional backgrounds revealed that the specific needs of their professions often shape users’ expectations of financial management apps. Business professionals sought tools to manage business-related finances efficiently, while doctors aimed to collect patient payments and related transactions effectively. These insights underscore the importance of tailoring financial apps to meet the unique requirements of diverse professions, ensuring effective financial management across various occupational backgrounds.
“As a doctor, it would be beneficial if I could easily keep track of all the patients who have paid their fees. Checking transaction histories repeatedly can be quite a hassle. So, having an automatically generated list of patients who have made payments, with a clear categorization of whether the payment has been completed, would be incredibly convenient.”-[P10]
Some respondents went beyond traditional budgeting and desired financial education and guidance within the app, along with receiving notifications and alerts when approaching spending limits or exceeding budgeted amounts.
“Besides handling expenses, it would be great if the app also offers guidance on achieving financial stability and discipline by providing tips and educational content to help users make better financial decisions.”-[P11]
@@ 5. Discussion
Answering RQ1 as mentioned in Section 1, the study results reveal a significant impact of UPI on Indian users’ spending habits. Most respondents (74.2%) accepted that UPI adoption led to increased spending, highlighting the UPI influence on user behavior. The data also indicates that many respondents started using UPI during significant economic and social changes, such as demonetization and pandemic-related lockdowns. Furthermore, 59.8% of respondents acknowledged exceeding their budget or overspending due to UPI usage. The qualitative analysis supports these findings, indicating that UPI has contributed to increased spending for most users. The abstract nature of digital transactions leads to psychological shifts, contributing to impulsive spending. In some cases, regret and guilt hit people after they realize their spending by checking their bank balances or statements, compared to spending physical cash, highlighting the lack of tangibility associated with UPI payments.
Answering RQ2 as mentioned in Section 1, user satisfaction levels were high, with 91.5% of the survey respondents reporting satisfaction with UPI usage and 95.2% of the survey respondents found making payments via UPI convenient. The qualitative analysis underscores the convenience of UPI transactions, with users praising its speed, efficiency, and the elimination of the need for physical cash. The absence of processing fees, as opposed to credit and debit cards, is also highlighted as a positive factor contributing to user satisfaction. The research findings indicate that users perceive UPI as a secure and efficient payment method, eliminating concerns about carrying large sums of money. This positive perception of convenience and efficiency likely contributes to the widespread adoption and satisfaction with UPI, ultimately shaping users’ preferences for cashless transactions and influencing spending behavior.
@@
Answering RQ3 as mentioned in Section 1, the study identifies variations in UPI adoption across demographic groups, particularly in professions. Working professionals and business owners integrate UPI into financial routines seamlessly, while students use it for academic needs. Despite variations, common spending trends are observed across age groups and professions, showcasing UPI’s adaptability to diverse user requirements.
5.1. RECOMMENDATIONS AND DESIRED FEATURES IN UPI & FINANCE TRACKING APPLICATIONS
@@ 5.2. FINANCIAL APP ANALYSIS IN INDIA
Stakeholders, including app developers, government agencies, and users, stand to benefit from incorporating our suggestions into future financial tracking app developments. This comprehensive analysis addresses user needs and provides valuable insights for enhancing the financial management ecosystem, ultimately advancing effective and user-centric solutions.
FEATURES YEARLY SUBSCRIPTION COST PROS CONS SUPPORTING PLATFORMS
FinArt (fin, 2016) Automatically organizes transactional SMS into expenses, budget, and bill reminders, even when the app is closed. Rs 799 (9.61 USD) Data stored on Drive, Auto Tracking, spend analysis dashboards, and budget setting. Separates personal and business accounts for comprehensive financial management. Setup complexity, ads persist after subscription, and occasional difficulty in SMS tracking. ANDROID, PC, IOS
Money Manager (mm, 2020) The app efficiently handles dual accounts, ensuring separate tracking of income and expenses. It offers offline/cloud backup and Data Export functionalities. Free Diverse payment method profiles, encompassing cash and cards, support competitive analysis. Easily create new profiles based on specific needs. The application takes a long time if currencies are changed to optimize expenses and income. Most fees are non-editable. ANDROID, IOS
Monefy (my, 2016) The app is straightforward; the infographics make it vibrant and easy to use. Rs 180 (2.17 USD) Set bill reminders, manage multiple accounts and easily import data through Excel. Frequent prompts for premium upgrades ANDROID, PC ,IOS
Wallet (wal, 2014) Users can use it for different currency rates worldwide. The Group Sharing option is also available in the application Rs 700 (8.42 USD) Monitor group expenses in the totals tab, track recent costs in the activity tab, and official API support available. Accounts auto-disconnect; unable to modify the initial login currency. ANDROID, IOS
Splitwise (spl, 2011) App simplifies bill splitting , not for personal expense tracking. Form groups, manage shared bills, and track dues effortlessly. Rs 499 (6 USD) Clean UI. Manages multiple wallets, links bank accounts or Crypto wallets directly, and tracks transactions seamlessly (premium feature). Tracking and analyzing personal finances can be challenging. ANDROID, PC, IOS
Spendee (spe, 2013) The app provides user-friendly expense and income tracking with easy-to-use tags for sorting costs efficiently. Rs 900 (10.43 USD) Effortlessly track expenses, share budgets, view spending analytics, and easily monitor debt progress. Certain features are restricted behind a paywall. ANDROID, PC, IOS
GoodBudget (gb, 2009) Budget management application. The app lets you create and refill multiple budgets (envelopes) at intervals like monthly or weekly. Rs 5817 (70 USD) Create multiple wallets, set up recurring payments, track credit card and debt. Link bank accounts and enable group sharing (premium only). Simple UI. Adding transactions is tedious, and there’s no API support. ANDROID, PC, IOS
Table 1. Comparison of Personal Finance Tracking Apps